# VLA-Pruner: Temporal-Aware Dual-Level Visual Token Pruning for Efficient Vision-Language-Action Inference

### v2

@@ 6 Conclusion
We propose VLA-Pruner, a training-free token pruning method for efficient VLA inference. To align with the VLA models’ dual-system nature of high-level semantic understanding and low-level action execution, VLA-Pruner employs a dual-level token importance criterion that integrates semantic-level relevance and action-level importance. Based on this criterion, our dual-level token selection strategy adaptively retains visual tokens. Across diverse robotic manipulation tasks and multiple VLA architectures, VLA-Pruner reduces computation while preserving performance. We discuss our limitations in Appendix C.
Supplementary Material
@@ A.1 Algorithm Outlines
We detail VLA-Pruner in Algorithm 1 (min-redundancy filtering) and Algorithm 2 (overall dual-level token selection).
@@ A.2 Complexity Analysis
Following FastV [8], SparseVLM [60] and DivPrune [2], we estimate the computational requirement of VLA-Pruner in terms of Transformer FLOPs, and focus on how pruning visual tokens changes the effective sequence length seen by each layer. As in prior work [2], we count the FLOPs of multi-head attention and feed-forward networks (FFNs) in the backbone (the overhead of token selection as negligible compared to the main model FLOPs.) Let be the total number of Transformer layers in the VLA backbone and the index of the pruning layer (we prune tokens before the layer). Denote the numbers of text and visual tokens before pruning by and (), and let be the full sequence length. After VLA-Pruner keeps a fraction of the visual tokens, the sequence length becomes . Following [8], the FLOPs of attention and FFN for one layer with sequence length , hidden size , and FFN intermediate size can be approximated as
@@
Since the attention term in is quadratic in sequence length, the reduction scales roughly with in layers after . This matches the empirical FLOPs ratios in Tab. 4.2: VLA-Pruner reduces FLOPs to about , , and of the original model at , , and token retention, respectively. Temporal smoothing over a short window of action-to-vision attention has complexity per timestep, and the greedy Max–Min filtering over the dual-level candidate pool of size is per prefill. In practice, these components contribute little to the total wall-clock time, as these operations are invoked only once during the prefill stage. Finally, Eq. (13) shows that VLA-Pruner offers a flexible way to trade accuracy for efficiency: the retained-token ratio . Smaller reduces FLOPs quadratically in layers after , yielding higher speedup, while larger recovers the original model’s performance. In our experiments, we report results at , but VLA-Pruner can be used at any intermediate ratio to meet a given compute or latency target.
@@ B.1.1 VLA Model Details
We consider three representative VLA architectures in our experiments: OpenVLA, OpenVLA-OFT, and . All models take RGB observations and a natural language instruction as input, and produce robot actions end-to-end. OpenVLA. OpenVLA [22] is a 7B-parameter open-source vision–language–action model built by fine-tuning a Prismatic VLM (Llama 2 language backbone with DINOv2 and SigLIP visual encoders) on robot manipulation episodes from the Open X-Embodiment dataset. The model represents actions as discrete tokens and decodes them autoregressively, enabling a single policy to control multiple embodiments out-of-the-box. OpenVLA serves as our base autoregressive VLA, and we follow its official checkpoints and evaluation settings on LIBERO. OpenVLA-OFT. OpenVLA-OFT [24] is an “Optimized Fine-Tuning” variant of OpenVLA that improves both control performance and inference efficiency. It augments the original decoder with parallel decoding and action chunking, and switches to a continuous action representation trained with a simple L1 regression objective. This substantially boosts success rates on LIBERO while increasing action-generation throughput. To be noted, OpenVLA-OFT utilizes bidirectional attention, meaning that during inference, there is already an action-to-vision cross-attention. However, we observe that during the early-layer prefill phase, the attention still focuses on semantic understanding, making action-to-vision less significant. In deeper layers, the action-to-vision attention starts to exhibit more distinct properties. This allows our method to effectively enhance token pruning performance by utilizing past timestep action-to-vision cross-attention at the deeper layers to guide current token pruning. Model. [4] is a prototype generalist robot policy that augments a pretrained vision–language model (PaliGemma backbone) with a continuous-action flow-matching head. Built on top of a large VLM to inherit Internet-scale semantic knowledge, is trained on diverse real-robot data and is designed for high-frequency control in dexterous manipulation tasks. The model decodes continuous motor commands via a flow-matching process with 10 integration steps. We use as a diffusion-based VLA backbone to evaluate the cross-architecture generalizability of VLA-Pruner. To capture informative interactions and reduce noise, we record the action-to-vision cross-attention scores averaged over the latter half of these integration steps.
@@ B.1.2 Acceleration Baseline Details
FastV. FastV [8](ECCV24) accelerates large vision–language models by pruning redundant visual tokens based on early-layer prefill attention. Its core observation is that many image patches receive consistently low attention in deeper layers while contributing substantially to computation. FastV defines an attention-based importance score during the prefill stage and drops a fixed fraction of low-score visual tokens after a chosen layer. It is training-free and purely semantic, but does not account for the action-dependent requirements of VLA models.
@@
SparseVLM. SparseVLM [60](ICML25) also targets visual token redundancy in VLMs, but uses text-to-vision cross-attention as the importance signal. Tokens that receive low cross-attention from text queries are considered less relevant and pruned. Compared to FastV, SparseVLM focuses directly on aligning visual tokens with language semantics, yet still operates only at the semantic level and ignores the distinct visual needs of action decoding in VLA.
@@
DivPrune. DivPrune [2](CVPR25) addresses the tendency of attention-based methods to select similar and redundant tokens. Instead of ranking tokens by attention scores, it formulates token selection as a Max–Min Diversity Problem and chooses a subset of visual tokens that are maximally dissimilar in feature space, thereby reducing redundancy while maintaining coverage. However, DivPrune is agnostic to the dual-level relevance of visual tokens in VLA prefilling and action decoding, and its diversity objective falls short for VLA tasks where localized details are critical.
@@
VLA-Cache. VLA-Cache [55](Neurips25) is a training-free acceleration method tailored for VLA models. It exploits temporal continuity in robotic manipulation by identifying visually static tokens whose features change little across timesteps, and reuses their computation via a KV-cache. At the same time, it uses text-to-vision cross-attention to keep task-relevant tokens fully computed. This cache-based mechanism reduces repeated computation over static background regions, but its semantic attention score based token selection remains coarse-grained, which overlooks distinct action attention distribution. To ensure a fair comparison, we set the dynamic token limit to be ( is the token budget), and use text-to-vision score to exclude of the most relevant tokens. These settings follow the default setup of VLA-Cache codebase.
@@
EfficientVLA. EfficientVLA [56](Neurips25) proposes a structured acceleration framework for diffusion-based VLA models that combines visual token pruning, layer reduction, and diffusion temporal feature caching. Its token-pruning component leverages semantic attention and diversity heuristics inherited from VLM pruning, while additional modules (i.e., layer skipping and diffusion caching) are used to achieve comparable acceleration. In contrast, VLA-Pruner is a single, plug-and-play token pruning module. These approaches are therefore complementary: EfficientVLA and other VLA acceleration methods can be adopted orthogonally with VLA-Pruner, while VLA-Pruner achieves strongest token-prune performance.
@@ B.2 Benchmark Details
LIBERO [32] is a Robosuite-based benchmark for lifelong robot manipulation that groups 130 language-conditioned tasks into four suites: Spatial, Object, Goal, and Long. LIBERO-Spatial varies only the spatial configuration of otherwise identical objects (e.g., pick up the black bowl on the stove and place it on the plate). LIBERO-Object varies the manipulated object while keeping the scene layout and goal template fixed (e.g., pick up the bbq sauce and place it in the basket). LIBERO-Goal keeps both the objects and spatial layout fixed but changes the goal description (e.g., open the middle drawer of the cabinet). LIBERO-Long corresponds to the long-horizon subset (10 tasks) of LIBERO-100, where each episode consists of multi-stage instructions such as opening a drawer, inserting an object, and then rearranging other items(e.g., put both the alphabet soup and the cream cheese box in the basket).
@@
SIMPLER [27] is a simulation suite built to closely track the real Google robot environment while allowing controlled appearance variation. It provides two evaluation settings. The Visual Matching (VM) setting keeps backgrounds, lighting, distractors, and camera poses close to the real setup so that rendered observations visually resemble real robot scenes and better predict real-world performance. The Variant Aggregation (VA) setting starts from the same base scenes but procedurally perturbs nuisance factors such as background textures, illumination, number and placement of distractor objects, table material, and camera viewpoints, creating a broad distribution shift for testing robustness. For the Google robot configuration, SIMPLER defines four canonical manipulation tasks: “pick coke can”,“pick object”, “move near”, “open/close Drawer”, and “place in closed drawer”. In our experiments, we follow the official simulator settings and success metrics, and select three representative tasks—Pick Coke Can, Move Near, and Open/Close Drawer—under VM or VA setting, where base model attains non-trivial performance.

### v3

@@ 6 Conclusion
We propose VLA-Pruner, a training-free token pruning method for efficient VLA inference. We reveal the dual-system nature of VLA models, exposing fundamental limitations in current pruning approaches. Considering this, VLA-Pruner employs a dual-objective importance criterion that integrates semantic-level relevance and action-level importance. Based on this, our proposed dual-level token selection strategy adaptively retains visual tokens. Across diverse robotic manipulation tasks and multiple VLA architectures, VLA-Pruner reduces computation while preserving performance. We discuss our limitations in Appendix C.
@@ A.2 Complexity Analysis
Following FastV (Chen et al., 2024), SparseVLM (Zhang et al., 2024) and DivPrune (Alvar et al., 2025), we estimate the computational requirement of VLA-Pruner in terms of Transformer FLOPs, and focus on how pruning visual tokens changes the effective sequence length seen by each layer. As in prior work (Alvar et al., 2025), we count the FLOPs of multi-head attention and feed-forward networks (FFNs) in the backbone (the overhead of token selection as negligible compared to the main model FLOPs.) Let be the total number of Transformer layers in the VLA backbone and the index of the pruning layer (we prune tokens before the layer). Denote the numbers of text and visual tokens before pruning by and (), and let be the full sequence length. After VLA-Pruner keeps a fraction of the visual tokens, the sequence length becomes . Following (Chen et al., 2024), the FLOPs of attention and FFN for one layer with sequence length , hidden size , and FFN intermediate size can be approximated as
@@
Since the attention term in is quadratic in sequence length, the reduction scales roughly with in layers after . This matches the empirical FLOPs ratios in Tab. 5: VLA-Pruner reduces FLOPs to about , , and of the original model at , , and token retention, respectively. Temporal smoothing over a short window of action-to-vision attention has complexity per timestep, and the greedy Max–Min filtering over the dual-level candidate pool of size is per prefill. In practice, these components contribute little to the total wall-clock time, as these operations are invoked only once during the prefill stage. Finally, Eq. (13) shows that VLA-Pruner offers a flexible way to trade accuracy for efficiency: the retained-token ratio . Smaller reduces FLOPs quadratically in layers after , yielding higher speedup, while larger recovers the original model’s performance. In our experiments, we report results at , but VLA-Pruner can be used at any intermediate ratio to meet a given compute or latency target.
@@ B.1.1 VLA Model Details
We consider three representative VLA architectures in our experiments: OpenVLA, OpenVLA-OFT, and . All models take RGB observations and a natural language instruction as input, and produce robot actions end-to-end. OpenVLA. OpenVLA (Kim et al., 2024) is a 7B-parameter open-source vision–language–action model built by fine-tuning a Prismatic VLM (Llama 2 language backbone with DINOv2 and SigLIP visual encoders) on robot manipulation episodes from the Open X-Embodiment dataset. The model represents actions as discrete tokens and decodes them autoregressively, enabling a single policy to control multiple embodiments out-of-the-box. OpenVLA serves as our base autoregressive VLA, and we follow its official checkpoints and evaluation settings on LIBERO. OpenVLA-OFT. OpenVLA-OFT (Kim et al., 2025b) is an “Optimized Fine-Tuning” variant of OpenVLA that improves both control performance and inference efficiency. It augments the original decoder with parallel decoding and action chunking, and switches to a continuous action representation trained with a simple L1 regression objective. This substantially boosts success rates on LIBERO while increasing action-generation throughput. To be noted, OpenVLA-OFT utilizes bidirectional attention, meaning that during inference, there is already an action-to-vision cross-attention. However, we observe that during the early-layer prefill phase, the attention still focuses on semantic understanding, making action-to-vision less significant. In deeper layers, the action-to-vision attention starts to exhibit more distinct properties. This allows our method to effectively enhance token pruning performance by utilizing past timestep action-to-vision cross-attention at the deeper layers to guide current token pruning. Model. (Black et al., 2024) is a prototype generalist robot policy that augments a pretrained vision–language model (PaliGemma backbone) with a continuous-action flow-matching head. Built on top of a large VLM to inherit Internet-scale semantic knowledge, is trained on diverse real-robot data and is designed for high-frequency control in dexterous manipulation tasks. The model decodes continuous motor commands via a flow-matching process with 10 integration steps. We use as a diffusion-based VLA backbone to evaluate the cross-architecture generalizability of VLA-Pruner. To capture informative interactions and reduce noise, we record the action-to-vision cross-attention scores averaged over the latter half of these integration steps.
@@ B.1.2 Acceleration Baseline Details
FastV. FastV (Chen et al., 2024)(ECCV24) accelerates large vision–language models by pruning redundant visual tokens based on early-layer prefill attention. Its core observation is that many image patches receive consistently low attention in deeper layers while contributing substantially to computation. FastV defines an attention-based importance score during the prefill stage and drops a fixed fraction of low-score visual tokens after a chosen layer. It is training-free and purely semantic, but does not account for the action-dependent requirements of VLA models.
@@
SparseVLM. SparseVLM (Zhang et al., 2024)(ICML25) also targets visual token redundancy in VLMs, but uses text-to-vision cross-attention as the importance signal. Tokens that receive low cross-attention from text queries are considered less relevant and pruned. Compared to FastV, SparseVLM focuses directly on aligning visual tokens with language semantics, yet still operates only at the semantic level and ignores the distinct visual needs of action decoding in VLA.
@@
DivPrune. DivPrune (Alvar et al., 2025)(CVPR25) addresses the tendency of attention-based methods to select similar and redundant tokens. Instead of ranking tokens by attention scores, it formulates token selection as a Max–Min Diversity Problem and chooses a subset of visual tokens that are maximally dissimilar in feature space, thereby reducing redundancy while maintaining coverage. However, DivPrune is agnostic to the dual-level relevance of visual tokens in VLA prefilling and action decoding, and its diversity objective falls short for VLA tasks where localized details are critical.
@@
VLA-Cache. VLA-Cache (Xu et al., 2025)(Neurips25) is a training-free acceleration method tailored for VLA models. It exploits temporal continuity in robotic manipulation by identifying visually static tokens whose features change little across timesteps, and reuses their computation via a KV-cache. At the same time, it uses text-to-vision cross-attention to keep task-relevant tokens fully computed. This cache-based mechanism reduces repeated computation over static background regions, but its semantic attention score based token selection remains coarse-grained, which overlooks distinct action attention distribution. To ensure a fair comparison, we set the dynamic token limit to be ( is the token budget), and use text-to-vision score to exclude of the most relevant tokens. These settings follow the default setup of VLA-Cache codebase.
@@
EfficientVLA. EfficientVLA (Yang et al., 2025)(Neurips25) proposes a structured acceleration framework for diffusion-based VLA models that combines visual token pruning, layer reduction, and diffusion temporal feature caching. Its token-pruning component leverages semantic attention and diversity heuristics inherited from VLM pruning, while additional modules (i.e., layer skipping and diffusion caching) are used to achieve comparable acceleration. In contrast, VLA-Pruner is a single, plug-and-play token pruning module. These approaches are therefore complementary: EfficientVLA and other VLA acceleration methods can be adopted orthogonally with VLA-Pruner, while VLA-Pruner achieves strongest token-prune performance.
[…]
@@ B.2 Benchmark Details
LIBERO (Liu et al., 2023a) is a Robosuite-based benchmark for lifelong robot manipulation that groups 130 language-conditioned tasks into four suites: Spatial, Object, Goal, and Long. LIBERO-Spatial varies only the spatial configuration of otherwise identical objects (e.g., pick up the black bowl on the stove and place it on the plate). LIBERO-Object varies the manipulated object while keeping the scene layout and goal template fixed (e.g., pick up the bbq sauce and place it in the basket). LIBERO-Goal keeps both the objects and spatial layout fixed but changes the goal description (e.g., open the middle drawer of the cabinet). LIBERO-Long corresponds to the long-horizon subset (10 tasks) of LIBERO-100, where each episode consists of multi-stage instructions such as opening a drawer, inserting an object, and then rearranging other items(e.g., put both the alphabet soup and the cream cheese box in the basket).
@@
SIMPLER (Li et al., 2024b) is a simulation suite built to closely track the real Google robot environment while allowing controlled appearance variation. It provides two evaluation settings. The Visual Matching (VM) setting keeps backgrounds, lighting, distractors, and camera poses close to the real setup so that rendered observations visually resemble real robot scenes and better predict real-world performance. The Variant Aggregation (VA) setting starts from the same base scenes but procedurally perturbs nuisance factors such as background textures, illumination, number and placement of distractor objects, table material, and camera viewpoints, creating a broad distribution shift for testing robustness. For the Google robot configuration, SIMPLER defines four canonical manipulation tasks: “pick coke can”,“pick object”, “move near”, “open/close Drawer”, and “place in closed drawer”. In our experiments, we follow the official simulator settings and success metrics, and select three representative tasks—Pick Coke Can, Move Near, and Open/Close Drawer—under VM or VA setting, where base model attains non-trivial performance.